OpenStudy (anonymous):

In converting 14 centimeters to inches, what unit (omit the number) would you place in the numerator of your ratio?

OpenStudy (mathstudent55):

Look at the fraction \( \dfrac{a}{b} \) below: \( 14 ~cm \times \dfrac{a}{b} \) You need to multiply 14 cm by the conversion ratio, and you need the cm units to canccel out so all you have left is inches. Where do you put 2.54 cm (numerator or denominator), and where do you put 1 inch (numerator and denominator) in order to cancel the cm units and to be left with inches?